Operations Report – December 3rd

Crew 216 Operations Report 03 Dec-19
SOL: 9
Name of person filing report: Mike Lawson
Non-nominal systems: NA
Generator: run
Hours run: 11.6
From what time last night: 1925
To what time this morning: 0657
List any additional daytime hours when the generator was run: N/A Solar— SOC % (Before generator is run at night: 59 %
Diesel Reading – 35
Station Propane Reading – 53
Ethanol Free Gasoline: N/A
Water (loft tank) (gal): 47
Water Meter (units): 1464585
Water (static tank) (gal): 236
Static to Loft Pump used – yes
Water in Green Hab (gal): 251
Water in Science Dome (gal): 0
Toilet tank emptied: no
Deimos rover used: no
Beginning charge:
Ending charge:
Currently charging:
Sojourner rover used:
Beginning charge:
Ending charge:
Currently charging:
Spirit rover used: yes x 2
Hours: 115.8; 116
Beginning charge: (Before EVA): 100; 100
Ending charge: (On return from EVA, before recharging): 81%; 79% Currently charging:
Opportunity rover used: no
Hours: 68.7
Beginning charge: 100
Ending charge: 100
Currently charging: yes
Curiosity rover used: yes x 2
Hours: 120.8; 121
Beginning charge: 100; 100
Ending charge: (On return from EVA, before charging): 77%; 85% Currently charging: yes
Notes on rovers: Opportunity to be moved off-site for maintenance. ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): no
Reason for use: N/A
Oil Added? no
# Hours the ATVs were used today: 0
Notes on ATVs:
HabCar used and why, where?
CrewCar used and why, where?
General notes and comments: EVA to Kissing Camel Ridge, rehearsal and prep for afternoon low light
EVA. UV mineral survey sites identified. Low light EVA provided proof of concept, samples collected for further analysis.
Summary of internet: No serious issues.
Summary of suits and radios: No issues.
Summary of Hab operations: Routine maintenance and housekeeping tasks. New hot water tank installation caused an electrical problem with the generator; HW system shut down at the direction of the Director; system later repaired and switched to propane.
Summary of GreenHab operations: routine watering.
Summary of Science Dome operations: Gypsum processing near final stage. Summary of RAM operations: RAM Rover power charging stations inop, electrical issues remain unresolved.
Summary of any observatory issues: NA
Summary of health and safety issues: NA
Questions, concerns and requests to Mission Support: 2 bag NB, 1 bag B type trash in the rear airlock.

Copyright © The Mars Society. All rights reserved. | Main Site